# Strong-to-Weak Spontaneous Symmetry Breaking in Mixed Quantum States

### v1

@@ Abstract
Symmetry in mixed quantum states can manifest in two distinct forms: strong symmetry, where each individual pure state in the quantum ensemble is symmetric with the same charge, and weak symmetry, which applies only to the entire ensemble. This paper explores a novel type of spontaneous symmetry breaking (SSB) where a strong symmetry is broken to a weak one. While the SSB of a weak symmetry is measured by the long-ranged two-point correlation function $\mathrm{Tr}(O_{x}O^{\dagger}_{y}\rho)$, the strong-to-weak SSB (SW-SSB) is measured by the fidelity $F(\rho,O_{x}O^{\dagger}_{y}\rho O_{y}O^{\dagger}_{x})$, dubbed the fidelity correlator. We prove that SW-SSB is a universal property of mixed-state quantum phases, in the sense that the phenomenon of SW-SSB is robust against symmetric low-depth local quantum channels. We argue that a thermal state at a nonzero temperature in the canonical ensemble (with fixed symmetry charge) should have spontaneously broken strong symmetry. Additionally, we study non-thermal scenarios where decoherence induces SW-SSB, leading to phase transitions described by classical statistical models with bond randomness. In particular, the SW-SSB transition of a decohered Ising model can be viewed as the “ungauged” version of the celebrated toric code decodability transition. We confirm that, in the decohered Ising model, the SW-SSB transition defined by the fidelity correlator is the only physical transition in terms of channel recoverability. We also comment on other (inequivalent) definitions of SW-SSB, through correlation functions with higher Rényi indices.
@@ II Definition and Properties
II.1 Fidelity correlators
@@ II.2 Rényi-2 correlators
It is then straightforward to check that $F_{O}(x,y)\to O(1)$ while $R^{(2)}(x,y)\to 0$, namely the state $\rho$ has SW-SSB in the fidelity sense, but not in the Rényi-$2$ sense. Another example illustrating the inequivalence between the fidelity correlator and the Rényi-2 correlator will be provided in Section III, where these two correlators are mapped to observables in distinct statistical models.
@@ III.1 SW-SSB in thermal states
$F(i,j)$ $=\frac{1}{\cosh^{2}\beta}(1+\tanh^{N}\beta)^{-1}$ (51)
@@
Analytically continuing to $n=1/2$, $F^{(n)}$ becomes the fidelity correlator, which demonstrates our original proposal. In particular, Eq. (55) agrees with Eq. (51) with $\Delta=2$ being the excitation gap.
@@ III.2 $\mathbb{Z}_{2}$ SW-SSB in the quantum Ising model
which is exactly the partition function of the effective Hamiltonian $H_{\mathrm{eff}}$. In particular, if we take the limit $m,n\rightarrow 1/2$, the replicated fidelity (57) will be recovered back to Eq. (6), and effective Hamiltonian (61) becomes the Hamiltonian of random-bond Ising model (RBIM) along the Nishimori line Fan et al. (2023). Then the transition between a $\mathbb{Z}_{2}$ strongly symmetric state at $p<p_{c}$ and an SW-SSB state induced by a symmetric quantum channel at $p>p_{c}$ happens at $p_{c}\approx 0.109$. Our results are fully consistent with the decodability transition of the toric code Dennis et al. (2002); Wang et al. (2003); Bao et al. (2023); Fan et al. (2023); Chen and Grover (2024a), which can now be interpreted as the “gauged” version of the SW-SSB transition through Kramers-Wannier transform – recall that gauge symmetry is naturally strong since the total gauge charge is constrained to be trivial in the Hilbert space.
@@
Furthermore, this is another example of the inequivalence between two different definitions of SW-SSB through fidelity and Rényi-2 correlators. For this model, the long-range Rényi-2 correlator is a stronger condition than the long-range fidelity correlator, since $p_{c}<p_{c}^{(2)}$.
@@ III.3 U(1) SW-SSB in a quantum rotor model
To further substantiate the result, we also study the Rényi-2 transition of this model. We find that $\operatorname{Tr}\rho^{2}$ can be mapped to a variant of the Villain XY model. In fact, for small $\alpha$, the Boltzmann weight is well-approximated by that of the Villain XY model. We thus expect that there is a Rényi-2 $\mathrm{U}(1)$ SW-SSB phase for sufficiently small $\alpha$, and a transition at a finite value of $\alpha$ (details can be found in Appendix A).
@@ IV Recoverability
where $\xi$ defines the correlation length of the mixed state. The local recovery channel $\mathcal{R}_{\mathcal{E}}$ is constructed through the Petz recovery channel. Two observations are important for our subsequent discussion: (1) The Petz recovery channel preserves the strong symmetry of the mixed state if the symmetry is on-site, which can be verified by the explicit form of the recovery channel (see Appendix B).
(2) Building on this local recovery, a finite-depth circuit channel acting on the entire system can be recovered region by region, provided that at each intermediate step, the local gates being recovered can be recovered locally, i.e., Eq. (83) holds true for some chosen buffer region. The global recovery channel is also a circuit channel, with depth scaling logarithmically with system size $\log L$. We refer the readers to Appendix B for details regarding the recovery channel.
@@ V Summary and Outlook
    We also considered an alternative definition of SW-SSB, which appeared in previous literature Lee et al. (2023); Ma et al. (2023); Ma and Turzillo (2024), by the long-ranged Rényi-$2$ correlator Eq. (29) of charged local operators. The two definitions of SW-SSB are inequivalent: there are examples with long-range fidelity correlators but short-range Rényi-2 correlators. Moreover, there is no analogue of the stability theorems (Thm. 1 and 2) for the Rényi-$2$ correlator, as we showed through an explicit example in Sec. IV. So the SW-SSB defined through the Rényi-$2$ correlator is not a universal property of mixed-state quantum phases. However, a weaker statement holds: a state with SW-SSB in the Rényi-$2$ sense must still be nontrivial, in the sense that it is not symmetrically invertible (Thm. 3).
@@
    We proposed that SW-SSB is a generic feature of thermal states at nonzero temperature (Sec. III.1). In particular, we conjectured that if the canonical ensemble (i.e., with fixed charge sector) of a local symmetric Hamiltonian without a strong-to-nothing SSB, then it must have SW-SSB. We demonstrated the statement in certain commuting-projector models and provided plausibility arguments for general Hamiltonians.
@@ Acknowledgements.
We thank Yimu Bao, Tarun Grover, Timothy Hsieh, Zhu-Xi Luo, Max Metlitski, Shengqi Sang, Brian Swingle, Ruben Verresen, Xiao-Gang Wen, Yichen Xu, Carolyn Zhang, Zhehao Zhang, and Yijian Zou for inspiring discussions. RM is especially grateful to Shengqi Sang for many discussions and for explaining Ref Sang and Hsieh (2024). JHZ and MC are grateful to Hong-Hao Tu and Wei Tang for discussions and collaborations on a related project. RM was supported in part by Simons Investigator Award number 990660 and by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (No. 651440). JHZ and ZB are supported by a startup fund from the Pennsylvania State University. RM, JHZ, ZB, and CW are supported in part by the grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics. LAL acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through Discovery Grants. Research at Perimeter Institute (LAL and CW) is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. MC acknowledges support from NSF under award number DMR-1846109. Note added – We would like to draw the reader’s attention to a related work Sala et al. on SW-SSB to appear in the same arXiv listing.
Appendix A $\mathrm{U}(1)$ Rényi-2 SW-SSB in the rotor model

### v2

@@ Abstract
Symmetry in mixed quantum states can manifest in two distinct forms: strong symmetry, where each individual pure state in the quantum ensemble is symmetric with the same charge, and weak symmetry, which applies only to the entire ensemble. This paper explores a novel type of spontaneous symmetry breaking (SSB) where a strong symmetry is broken to a weak one. While the SSB of a weak symmetry is measured by the long-ranged two-point correlation function $\mathrm{Tr}(O_{x}O^{\dagger}_{y}\rho)$, the strong-to-weak SSB (SW-SSB) is measured by the fidelity $F(\rho,O_{x}O^{\dagger}_{y}\rho O_{y}O^{\dagger}_{x})$, dubbed the fidelity correlator. We prove that SW-SSB is a universal property of mixed-state quantum phases, in the sense that the phenomenon of SW-SSB is robust against symmetric low-depth local quantum channels. We also show that the symmetry breaking is “spontaneous” in the sense that the effect of a local symmetry-breaking measurement cannot be recovered locally. We argue that a thermal state at a nonzero temperature in the canonical ensemble (with fixed symmetry charge) should have spontaneously broken strong symmetry. Additionally, we study non-thermal scenarios where decoherence induces SW-SSB, leading to phase transitions described by classical statistical models with bond randomness. In particular, the SW-SSB transition of a decohered Ising model can be viewed as the “ungauged” version of the celebrated toric code decodability transition. We confirm that, in the decohered Ising model, the SW-SSB transition defined by the fidelity correlator is the only physical transition in terms of channel recoverability. We also comment on other (inequivalent) definitions of SW-SSB, through correlation functions with higher Rényi indices.
@@ II Definition and Properties
II.1 Fidelity correlator and stability theorem
@@ II.2 Rényi-2 correlator
It is then straightforward to check that $F_{O}(x,y)\to O(1)$ while $R^{(2)}(x,y)\to 0$, namely the state $\rho$ has SW-SSB in the fidelity sense, but not in the Rényi-$2$ sense. The opposite situation, in which $F_{O}(x,y)\to 0$ but $R^{(2)}(x,y)\to O(1)>0$, can also happen, as shown in Appendix A. Another example illustrating the inequivalence between the fidelity correlator and the Rényi-2 correlator will be provided in Section III, where these two correlators are mapped to observables in distinct statistical models.
@@ II.3 Local indistinguishability and detectability
It is straightforward to show that the above result on local detectability of explicit strong symmetry breaking for $\rho_{+}$ is also valid in the Rényi-$2$ sense.
@@ III.1 SW-SSB in thermal states
$F(i,j)$ $=\frac{1}{\cosh^{2}\beta}(1+\tanh^{N}\beta)^{-1}$
@@
Analytically continuing to $n=1/2$, $F^{(n)}$ becomes the fidelity correlator, which demonstrates our original proposal. In particular, Eq. (54) agrees with Eq. (III.1) with $\Delta=2$ being the excitation gap.
@@ III.2 $\mathbb{Z}_{2}$ SW-SSB in the quantum Ising model
which is exactly the partition function of the effective Hamiltonian $H_{\mathrm{eff}}$. In particular, if we take the limit $m,n\rightarrow 1/2$, the replicated fidelity (56) will be recovered back to Eq. (6), and effective Hamiltonian (60) becomes the Hamiltonian of random-bond Ising model (RBIM) along the Nishimori line Fan et al. (2023). Then the transition between a $\mathbb{Z}_{2}$ strongly symmetric state at $p<p_{c}$ and an SW-SSB state induced by a symmetric quantum channel at $p>p_{c}$ happens at $p_{c}\approx 0.109$. We note that the same result for the fidelity correlator can be established directly without using the replica trick, which is explained in Appendix B. Our results are fully consistent with the decodability transition of the toric code Dennis et al. (2002); Wang et al. (2003); Bao et al. (2023); Fan et al. (2023); Chen and Grover (2024a), which can now be interpreted as the “gauged” version of the SW-SSB transition through Kramers-Wannier transform – recall that gauge symmetry is naturally strong since the total gauge charge is constrained to be trivial in the Hilbert space.
@@
Furthermore, this is another example of the inequivalence between two different definitions of SW-SSB through fidelity and Rényi-2 correlators. For this model, the long-range Rényi-2 correlator is a stronger condition than the long-range fidelity correlator, since $p_{c}<p_{c}^{(2)}$. We note that it has been proposed in Ref. Lee et al. (2023) to use Rényi-$1$ quantity, such as the Fisher information or von Neumann relative entropy to define an “intrinsic” transition; based on its correspondence to the decodability transition in toric code Dennis et al. (2002), this transition was claimed to be the random-bond transition at $p_{c}$.
@@ III.3 U(1) SW-SSB in a quantum rotor model
To further substantiate the result, we also study the Rényi-2 transition of this model. We find that $\operatorname{Tr}\rho^{2}$ can be mapped to a variant of the Villain XY model. In fact, for small $\alpha$, the Boltzmann weight is well-approximated by that of the Villain XY model. We thus expect that there is a Rényi-2 $\mathrm{U}(1)$ SW-SSB phase for sufficiently small $\alpha$, and a transition at a finite value of $\alpha$ (details can be found in Appendix C).
@@ IV Recoverability
where $\xi$ defines the correlation length of the mixed state. The local recovery channel $\mathcal{R}_{\mathcal{E}}$ is constructed through the Petz recovery channel. Two observations are important for our subsequent discussion: (1) The Petz recovery channel preserves the strong symmetry of the mixed state if the symmetry is on-site, which can be verified by the explicit form of the recovery channel (see Appendix D).
(2) Building on this local recovery, a finite-depth circuit channel acting on the entire system can be recovered region by region, provided that at each intermediate step, the local gates being recovered can be recovered locally, i.e., Eq. (82) holds true for some chosen buffer region. The global recovery channel is also a circuit channel, with depth scaling logarithmically with system size $\log L$. We refer the readers to Appendix D for details regarding the recovery channel.
IV.1 Recoveribility within the SSB phase
@@
IV.2 Spontaneity and local irrecoverability
In the previous example, the CMI $I(A;C|B)$ in the SW-SSB phase is nonzero ($=\ln 2$) at large $l_{B}$. We now prove that such non-vanishing CMI is a universal feature of SW-SSB states, which is connected to the local non-recoverability of (symmetry-breaking) local channels.
Theorem 4.
If a mixed state $\rho$ has SW-SSB in the fidelity correlator sense Eq. (6), then $\exists$ some local channel $\mathcal{E}_{i}$ (supported around site $i$) that cannot be recovered locally, namely $\nexists$ local $\mathcal{E}^{\prime}_{i}$ such that $\mathcal{E}^{\prime}_{i}\circ\mathcal{E}_{i}[\rho]=\rho$.
Furthermore, for any local $\mathcal{E}^{\prime}_{i}$, the trace distance $D(\rho,\mathcal{E}^{\prime}_{i}\circ\mathcal{E}_{i}[\rho])=O(1)>0$. As a consequence, the CMI $I(A;C|B)_{\rho}$ (in the geometry of Fig. 2) must remain nonzero at large distance between $A$ and $C$, for a region $A$ containing the support of $\mathcal{E}_{i}$ around site $i$.
Proof.
Let us normalize the order parameter $O_{i}$ such that the following operation is a valid quantum channel:
$\mathcal{E}_{i}[\rho]\equiv\frac{1}{2}\rho+\frac{1}{2}O_{i}\rho O_{i}^{\dagger}+\frac{1}{2}\sum_{a}K_{i,a}\rho K_{i,a}^{\dagger},$ (89)
where the local Kraus operators $\{K_{i,a}\}$ are chosen so that $O_{i}^{\dagger}O_{i}+\sum_{a}K_{i,a}^{\dagger}K_{i,a}=I$. For example, if $O_{i}$ is unitary, then we can choose $K_{i}=0$. In the Ising example $O=Z$ and the channel is nothing but a strong $Z$-measurement.
Now consider another arbitrary channel around site $i$, call it $\mathcal{E}^{\prime}_{i}$, and define
$\tilde{\rho}\equiv\mathcal{E}^{\prime}_{i}\circ\mathcal{E}_{i}[\rho].$ (90)
To see that $\tilde{\rho}\neq\rho$, it suffices to show that for some site $j$ far away from site $i$, $F(\tilde{\rho},O_{j}^{\dagger}\tilde{\rho}O_{j})=O(1)>0$, namely $\tilde{\rho}$ does not even have the strong symmetry. This claim follows simply from the joint concavity and the data processing inequality of fidelity:
$F(\tilde{\rho},O_{j}^{\dagger}\tilde{\rho}O_{j})$$\geq$ $\frac{1}{2}F(\mathcal{E}^{\prime}_{i}[\rho],\mathcal{E}^{\prime}_{i}[O_{i}O_{j}^{\dagger}\rho O_{j}O_{i}^{\dagger}])$ (91)
$\geq$ $\frac{1}{2}F(\rho,O_{i}O_{j}^{\dagger}\rho O_{j}O_{i}^{\dagger})$
$=$ $O(1)>0.$
Next we examine more carefully the trace distance $D(\rho,\tilde{\rho})$. Using the triangle inequality of trace distance, we have
$D(\rho,\tilde{\rho})$$\geq$ $D(\rho,O_{j}^{\dagger}\rho O_{j})-D(\tilde{\rho},O_{j}^{\dagger}\rho O_{j})$
$\geq$ $D(\rho,O_{j}^{\dagger}\rho O_{j})-D(\tilde{\rho},O_{j}^{\dagger}\tilde{\rho}O_{j})-D(O_{j}^{\dagger}\rho O_{j},O_{j}^{\dagger}\tilde{\rho}O_{j}).$
Now we use several simple observations:
  1. (1)
    $D(\rho,O_{j}^{\dagger}\rho O_{j})=1$ by strong symmetry;
  2. (2)
    $D(\tilde{\rho},O_{j}^{\dagger}\tilde{\rho}O_{j})=1-c$ for some $c>0$ due to Eq. (91);
  3. (3)
    using the sub-multiplicative property of trace norm, we have
$D(O_{j}^{\dagger}\rho O_{j},O_{j}^{\dagger}\tilde{\rho}O_{j})\leq\|O_{j}\|^{2}D(\rho,\tilde{\rho}),$ (93)
where $\|O_{j}\|\sim O(1)$ is the operator norm of $O_{j}$. Putting all these relations back into Eq. (LABEL:eq:recovertracedistance), we obtain
$D(\rho,\tilde{\rho})\geq c/(1+\|O_{j}\|^{2})>0.$ (94)
The above relation holds for any attempted recovery channel $\mathcal{E}^{\prime}$. In particular, if we choose $\mathcal{E}^{\prime}_{i}$ to be the Petz recovery map $\mathcal{R}_{i}$, then we have (see Eq. (122) in Appendix D)
$D(\rho,\mathcal{R}_{i}\circ\mathcal{E}_{i}[\rho])\leq\sqrt{\ln 2\cdot I(A;C|B)_{\rho}},$ (95)
which immediately implies that the CMI $I(A;C|B)_{\rho}$ must remain nonzero at large ${\rm dist}(A,C)$. ∎
[…]
The above notion of spontaneity is “weak” in the sense that the irrecoverability holds for certain measurement results. If we average over (or simply forget) the measurement results, the post-measurement state may become locally recoverable. For example, take the weak-to-trivial example $\rho=\ket{\uparrow\cdots\uparrow}\!\bra{\uparrow\cdots\uparrow}+\ket{\downarrow\cdots\downarrow}\!\bra{\downarrow\cdots\downarrow}$ |, the post-measurement density matrix is $\mathcal{E}[\rho]=(1/2)\rho+(1/2)Z_{i}\rho Z_{i}=\rho$. If the post-measurement density matrix (without post-selecting the measurement outcome) cannot be locally recovered to the original state, we dub the spontaneity “strong” – this is a stronger statement since if the post-measurement state is locally recoverable for all measurement results, it is simple to construct a recovery channel that recovers the whole density matrix, by taking the convex sum of the corresponding recovery channel for each measurement result.
In this language, what we have shown in Thm. 4 is that the SSB of strong symmetry (either to weak or to trivial) is strongly spontaneous. Let us consider a pedagogical example of SW-SSB: $\rho_{0}=\mathbbm{1}+X$ as the SW-SSB of $\mathbb{Z}_{2}$ Ising symmetry (Sec. II.1). We perform a strong $Z$-measurement on a single qubit and recognize that $\rho_{0}$ will immediately collapse to the maximally mixed state $\mathbbm{1}$ that no longer has the strong Ising symmetry. Heuristically, the only information encoded in the SW-SSB density matrix is the global charge sector fixing, i.e., $\rho_{0}$ is the maximally mixed state within the even charge sector of the Ising symmetry. Once we perform a strong $Z$-measurement, the information on global charge sector fixing is completely destroyed, and we can never restore this information with any local quantum channel.
[…]
@@ V Summary and Outlook
    We also considered an alternative definition of SW-SSB, which appeared in previous literature Lee et al. (2023); Ma et al. (2023); Ma and Turzillo (2024), by the long-ranged Rényi-$2$ correlator Eq. (29) of charged local operators. The two definitions of SW-SSB are inequivalent: there are examples with long-range fidelity correlators but short-range Rényi-2 correlators, and vice-versa (Appendix A). Moreover, there is no analogue of the stability theorems (Thm. 1 and 2) for the Rényi-$2$ correlator, as we showed through an explicit example in Sec. IV. So the SW-SSB defined through the Rényi-$2$ correlator is not a universal property of mixed-state quantum phases. However, a weaker statement holds: a state with SW-SSB in the Rényi-$2$ sense must still be nontrivial, in the sense that it is not symmetrically invertible (Thm. 3).
@@
    We proposed that SW-SSB is a generic feature of thermal states at nonzero temperature (Sec. III.1). In particular, we conjectured that if the canonical ensemble (i.e., with fixed charge sector) of a local symmetric Hamiltonian without a strong-to-trivial SSB, then it must have SW-SSB. We demonstrated the statement in certain commuting-projector models and provided plausibility arguments for general Hamiltonians.
@@
  6. 6.
    We demonstrated the spontaneity of SW-SSB and its relation to the local irrecoverability of local symmetry-breaking channels. Specifically, we proved that for a density matrix with SW-SSB, a local quantum channel can induce a global change that cannot be locally recovered (Thm. 4). This further implied that a non-vanishing CMI is a universal feature of SW-SSB.
@@ Acknowledgements.
We thank Yimu Bao, Tarun Grover, Timothy Hsieh, Andy Lucas, Zhu-Xi Luo, Max Metlitski, Shengqi Sang, Brian Swingle, Ruben Verresen, Xiao-Gang Wen, Yichen Xu, Carolyn Zhang, Zhehao Zhang, and Yijian Zou for inspiring discussions. RM is especially grateful to Shengqi Sang for many discussions and for explaining Ref. Sang and Hsieh (2024). JHZ and MC are grateful to Hong-Hao Tu and Wei Tang for discussions and collaborations on a related project. RM was supported in part by Simons Investigator Award number 990660 and by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (No. 651440). JHZ and ZB acknowledge support from NSF under award number DMR-2339319. RM, JHZ, ZB, and CW are supported in part by the grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics. LAL acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through Discovery Grants. Research at Perimeter Institute (LAL and CW) is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. MC acknowledges support from NSF under award number DMR-1846109. Note added – We would like to draw the reader’s attention to a related work Sala et al. on SW-SSB to appear in the same arXiv listing.
Appendix A Counterexample for SW-SSB only in the Rényi-2 sense.
Here, we are going to show $N$-qubit states $\rho_{N}$ that are strongly symmetric under $\mathbb{Z}_{2}$, exhibit SW-SSB in the sense of Rényi-2 correlator (Definition 4), but not in the fidelity sense (Definition 1). The strategy will be to construct “spin glass” ensembles $\rho_{N}\propto\sum_{m=1}^{M_{N}}p_{m}\ket{\psi_{m}}\!\bra{\psi_{m}}$ |, $\lim_{N\to\infty}M_{N}=\infty$, such that $\lim_{|x-y|\to\infty}\lim_{N\to\infty}F_{O,N}(x,y)=0$ for all order parameters $O$, but with its Rényi-2 correlator satisfying
$\lim_{|x-y|\to\infty}\lim_{N\to\infty}R^{(2)}_{O^{\prime},N}(x,y)$ $=\frac{\sum_{m}p_{m}^{2}|o^{\prime}_{m}|^{2}}{\sum_{m}p_{m}^{2}}>0,$ (96)
for the particular order parameter $O^{\prime}=Z$, with $\lim_{|x-y|\to\infty}\braket{\psi_{m}}{Z_{x}Z_{y}}{\psi_{n}}=o^{\prime}_{m}\delta_{mn}$. Furthermore, we require $\sum_{m}p_{m}$ to diverge but $\sum_{m}p_{m}|o^{\prime}_{m}|,\sum_{m}p_{m}^{2}|o^{\prime}_{m}|^{2}$ and $\sum_{m}p_{m}^{2}$ to converge. In that way, at least for $o^{\prime}$, the fidelity correlator goes to zero, since
$\lim_{|x-y|\to\infty}\lim_{N\to\infty}F_{O^{\prime},N}(x,y)=\frac{\sum_{m}p_{m}|o^{\prime}_{m}|}{\sum_{m}p_{m}}=0,$ (97)
and we have dropped the normalization condition $\sum_{m}p_{m}=1$ for simplicity of calculation.
Let us consider a qubit chain with $2N$ spins, and define the strong $\mathbb{Z}_{2}$ symmetry to only act on even spins, meaning the symmetry is generated by $X_{e}:=\otimes_{n}X_{2n}$. We now consider all length-$N$ bit strings $m_{1}m_{2}\cdots m_{N}$, and view each bit string as the binary representation of an integer $m\in\{1,2,\ldots,2^{N}\}$. We can then construct the density matrix $\rho_{N}=\sum_{m=1}^{2^{N}}p_{m}\ket{\psi_{m}}\!\bra{\psi_{m}}$ |, where $p_{m}=m^{-2/3}$ and
$|\psi_{m}\rangle$ $=$ $\frac{\mathbbm{1}+X_{e}}{\sqrt{2}}|m_{1}\tilde{m}_{1}m_{2}\tilde{m}_{2}...\rangle,$
$|\tilde{m}_{i}\rangle:$ $=$ $\sqrt{\frac{1}{2}+\frac{1}{m}}|m_{i}\rangle+\sqrt{\frac{1}{2}-\frac{1}{m}}X_{i}|m_{i}\rangle,$ (98)
with $|m_{i}\rangle$ defined in the $Z$-basis. At large $m$, $|\tilde{m}_{i}\rangle\approx|+\rangle+[(-1)^{m_{i}}/m]|-\rangle$, so we should have $o^{\prime}_{m}\sim 1/m^{2}$. The auxiliary spins on the odd sites are introduced to make sure that different $|\psi_{m}\rangle$’s are orthogonal w.r.t. $O^{\prime}$, allowing the simple forms of the correlators as in Eq. (96).
Given that $p_{m}\sim m^{-2/3}$ and $o^{\prime}_{m}\sim 1/m^{2}$, then $\sum_{m}p_{m}|o^{\prime}_{m}|,\sum_{m}p_{m}^{2}|o^{\prime}_{m}|^{2}$ and $\sum_{m}p_{m}^{2}$ all converge to an $O(1)$ value, while $\sum_{m}p_{m}\sim m_{\rm max}^{1/3}\sim 2^{N/3}$. We conclude that while the Rényi-2 remains finite, the fidelity correlator for $O^{\prime}=Z$ vanishes exponentially in the thermodynamic limit. For other order parameters, the fidelity correlator still goes to zero, as $\rho_{N}$ converges to the trivial state $\rho_{\rm odd}\otimes\ket{+}\!\bra{+}_{\rm even}$, $\rho_{\rm odd}\propto\sum_{m}p_{m}\ket{m}\!\bra{m}$ |, in the fidelity (or trace) distance as $N\to\infty$:
$F(\rho_{N},\rho_{\rm odd}\otimes\ket{+}\!\bra{+}_{\rm even})\approx\frac{\sum_{m}p_{m}(1-N/2m^{2})}{\sum_{m}p_{m}}\to 1.$ (99)
It is clear from this construction, however, that this sequence of states is highly artificial. Thus, it remains a question of whether the same can happen for states arising from more natural contexts, such as thermal or decohered states.
Appendix B Alternative derivation of the fidelity correlator
We present an alternative derivation of the fidelity correlator in the decohered 2D Ising model Eq. (55) without using the replica trick.
The starting point is the decohered density matrix in the $\mathbb{Z}_{2}$ charge basis:
$\rho=\sum_{\braket{X}}Z(X)\ket{X}\bra{X}.$ | . (100)
Here $X$ denotes a configuration of $\mathbb{Z}_{2}$ charges (i.e. $X=\pm 1$), subject to the constraint that $\prod_{i}X_{i}=1$. In the following we write $X_{i}=(-1)^{n_{i}}$ with $n_{i}=0,1$. To compute the weight $Z(X)$, define an edge $\mathbb{Z}_{2}$ variable $s_{ij}=0,1$, which keeps track of the Kraus operator on the edge $ij$. When applied to the initial state |$\ket{X_{i}=1}$, the resulting charge at site $i$ is given by
$(\nabla\cdot s)_{i}=\sum_{\mu=x,y}s_{i,i+\mathbf{e}_{\mu}}+s_{i-\mathbf{e}_{\mu},i}\text{ mod }2.$ (101)
Here $\mathbf{e}_{\mu}$ is the unit vector along the $\mu$ direction. We obtain
$Z(X)$ $=\sum_{\nabla\cdot s=n}\prod_{\langle ij\rangle}(1-p)^{1-s_{ij}}p^{s_{ij}}$
$=(1-p)^{N_{e}}\sum_{\nabla\cdot s=n}\Big{(}\frac{p}{1-p}\Big{)}^{\sum_{\langle ij\rangle}s_{ij}}.$ (102)
To resolve the constraint $\nabla\cdot s=n$, we denote a special solution of this equation by $s^{0}$, and $s_{e}=s^{0}_{e}+\tilde{s}_{p}+\tilde{s}_{q}$ mod 2 (where $e$ stands for a nearest-neighbor edge, and $p,q$ are the two adjacent square plaquettes). $\tilde{s}_{p}$ are dual variables living on the plaquettes $p$. It is more convenient to work with $\mu_{e}=(-1)^{s_{e}}$ and $\tau_{p}=(-1)^{\tilde{s}_{p}}$, so $\mu_{e}=\mu_{e}(X)\tau_{p}\tau_{q}$, with $\mu_{e}(X)=(-1)^{s_{e}^{0}}$. Therefore
$Z(X)=[p(1-p)]^{N_{e}/2}\sum_{\tau}e^{\beta\sum_{\langle pq\rangle}\mu_{pq}(X)\tau_{p}\tau_{q}}$ (103)
where $e^{-2\beta}=\frac{p}{1-p}$. In other words, $Z(X)$ is the partition function of a 2D classical Ising model coupled to background $\mathbb{Z}_{2}$ gauge fields $\{\mu(X)\}$, with fixed gauge-invariant $\mathbb{Z}_{2}$ fluxes given by $X$.
Since $\rho$ is already diagonalized in this basis, we can directly compute the fidelity correlator. We have
$\sigma$ $=Z_{i}Z_{j}\rho Z_{i}Z_{j}$
$=\sum_{X}Z(X\cdot\delta X)\ket{X}\bra{X}.$ | . (104)
where $\delta X_{k}=-1$ if $k=i,j$ and 1 otherwise, and the fidelity correlator is found to be Chen and Grover (2024a)
$F_{ij}$ $=\operatorname{Tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}$ (105)
$=\sum_{X}\sqrt{Z(X)Z(X\cdot\delta X)}$
$=\sum_{X}Z(X)\left(\frac{Z(X\cdot\delta X)}{Z(X)}\right)^{1/2}$
$=\sum_{X}Z(X)e^{-\frac{1}{2}\Delta F_{\mathrm{d}w}}$
$=\braket{e^{-\frac{1}{2}\Delta F_{\mathrm{d}w}}}_{\rm RBIM}\geq e^{-\frac{1}{2}\braket{\Delta F_{\rm dw}}_{\rm RBIM}}.$ (106)
Here $\Delta F_{\rm dw}=-\ln\frac{Z(X\cdot\delta X)}{Z(X)}$ is the free energy cost of inserting a domain wall of length $\sim|i-j|$. $\braket{\cdots}_{\rm RBIM}$ stands for disorder average in the random-bond Ising model along the Nishimori line Dennis et al. (2002). For $p>p_{c}$, the RBIM is in the paramagnetic phase (i.e. high temperature), so the disorder-averaged free energy cost for a domain wall approaches a constant. Thus we have shown that the fidelity correlator is long-range ordered for $p>p_{c}$. It also suggests that $F_{ij}$ decays exponentially for $p<p_{c}$, since the free energy cost diverges linearly with the distance in the ferromagnetic (low-temperature) phase. We note that the same results follow for the other definition of SSB in terms of sandwiched Rényi entropy.
Appendix C $\mathrm{U}(1)$ Rényi-2 SW-SSB in the rotor model